OpenStudy (anonymous):

An airplane flying at an altitude of 32,000 feet descends at a rate of 1,300 feet per minute. If the plane descends for 15 minutes, what would its altitude be in feet?

OpenStudy (anonymous):

if the angle of descending is not given then:  first u need to find the distance traveled during the 15min. to do so: v=d/t so d=v.t=(1300ft/min).15min=19500ft. but initially the plane is at 32000ft altitude, from there it travelled 19500ft, so the altitude(relative to ground position,0,) of the final position is 32000-19500=12500ft as illustrated in drawing hope that helps